# Kernel-based Koopman approximants for control: Flexible sampling, error analysis, and stability

### v1

@@ 4. Koopman approximants for control systems
where $x\in\Omega\subset\mathbb{R}^{n}$ and the control $u$ is contained in a bounded set $\mathbb{U}\subset\mathbb{R}^{m}$, $m\leq n$, with $\operatorname{span}\mathbb{U}=\mathbb{R}^{m}$. Here, we assume that the maps $g_{0}:\Omega\to\mathbb{R}^{n}$ and $G:\Omega\to\mathbb{R}^{n\times m}$ are locally Lipschitz continuous, i.e., there exist constants $L_{g}>0$ and $L_{G}>0$ such that
$\|g_{0}(x)-g_{0}(y)\|\leq L_{g}\|x-y\|\qquad\text{and}\qquad\|G(x)-G(y)\|\leq L_{G}\|x-y\|$
for all $x,y\in\Omega$. We furthermore assume that all entries of both $g_{0}$ and $G$ are functions in the RKHS $\mathbb{H}=\mathcal{N}_{\Phi_{n,k}}(\Omega)$ for some fixed $k\in\mathbb{N}$. Hereby, we note that, in view of Equation 3, $\mathcal{N}_{\Phi_{n,k_{1}}}(\Omega)\subset\mathcal{N}_{\Phi_{n,k_{2}}}(\Omega)$ for $k_{1}\geq k_{2}$.
@@
Often, systems Equation 23 are derived from continuous-time control-affine systems governed by the dynamics
$\dot{x}(t):=\tilde{f}(x(t),u(t)):=\tilde{g}_{0}(x(t))+\sum_{i=1}^{m}u_{i}(t)\tilde{g}_{i}(x(t))$ (24)
with locally Lipschitz-continuous maps $\tilde{g}_{i}:\Omega\rightarrow\mathbb{R}^{n}$, $i\in[0:m]$. Analogously to the autonomous case, see Remark 2.1, for a time step $\Delta t>0$ we obtain
$x(\Delta t;\bar{x},u)=\bar{x}+\int_{0}^{\Delta t}\tilde{g}_{0}(x(s;\bar{x},u))\,\mathrm{d}s+\int_{0}^{\Delta t}\tilde{G}(x(s;\bar{x},u))u(s)\,\mathrm{d}s,$
where $\tilde{G}$ is defined analogously to $G$ above – again tacitly assuming existence and uniqueness of the solution $x(\cdot;\bar{x},u)$ on $[0,\Delta t]$. Then, assuming that the control function $u$ is constant on the interval $(0,\Delta t)$, a Taylor-series expansion of the solution at $t=0$ yields
$x(\Delta t;\bar{x},u)=\bar{x}+\Delta t\left[\tilde{g}_{0}(\bar{x})+\tilde{G}(\bar{x})u\right]+\mathcal{O}(\Delta t^{2}),$
where we have invoked compactness of $\mathbb{U}$ and Lipschitz continuity of $f(\cdot,u)$ and $\frac{\partial}{\partial x}f(\cdot,u)$ on $\overline{\Omega}$. Hereby, we have identified the (constant) control functions with the respective control value $u\in\mathbb{U}\subset\mathbb{R}^{m}$. Thus, we obtain a discrete-time system Equation 23 up to an arbitrarily small error if the time step $\Delta t$ is sufficiently small. This is in alignment with the theory of sampled-data systems with zero-order hold, see, e.g., [29].
In this section, we propose a novel learning architecture to learn Koopman approximants of the control-affine system Equation 23, that allows for flexible sampling of state-control pairs $(\bar{x}_{i},\bar{u}_{i},\bar{x}_{i}^{+})$ with $\bar{x}_{i}\in\Omega$, $\bar{u}_{i}\in\mathbb{U}$, and $\bar{x}_{i}^{+}=f(\bar{x}_{i},\bar{u}_{i})$, $i\in[1:d]$. The result of our proposed kEDMD-based algorithm are control-affine surrogate dynamics
@@
using regression and interpolation in Wendland native spaces. In Theorem 4.3, we prove a uniform bound on the approximation error $\|f(x,u)-\widehat{f}(x,u)\|$, similarly as in Corollary 2.8 for the approximants $\widehat{F}$ and $\widehat{F}_{\lambda}$ of the dynamical system Equation DS. Then, we provide an application of the derived proportional bound in view of feedback stabilization. To be more precise, we prove in Section 4.2 that if a feedback stabilizes the surrogate dynamics Equation 25, the desired set point is stabilized in the original system (and vice versa). Last, the established error bounds are validated numerically in Section 4.3.
@@ 4.1. Koopman approximants, flexible sampling, and error bounds
We propose a novel learning architecture to generate Koopman approximants for control-affine systems with the following two key features: On the one hand, the algorithm can be used with (almost) arbitrary state-control data allowing for highly flexible sampling including i.i.d., ergodic, and trajectory-based data generation. On the other hand, we rigorously show bounds on the full approximation error with explicit convergence rates in the infinite-data limit and, foremost, allow for controller design with closed-loop guarantees without imposing restrictive invariance assumptions on the dictionary in EDMD. To the best of our knowledge, this combination is, up to now, unique. We note that Koopman-based approximations of control-affine systems in an RKHS framework with flexible sampling were also presented in [3], where the kernel (and thus also the RKHS) is suitably extended to capture the control dependency. Therein, however, no error bounds were provided. In addition, our novel approach allows to counteract numerical ill-conditioning by decomposing the approximation process in two steps labeled as macro and micro level, see the following paragraph for details.
The proposed method provided in detail in Algorithm 1 consists of two steps. In a first step, we determine $d$ clusters of size $N\geq m+1$, where $m$ is the control dimension in Equation 23. Each cluster corresponds to a center $x_{\ell}\in\Omega$, $\ell\in[1:d]$, and its $N$ nearest neighbors. The chosen centers $x_{\ell}$, $\ell\in[1:d]$, determine the fill distance $h_{\mathcal{X}}$. Then, we use the data triples $(\bar{x}_{i},\bar{u}_{i},\bar{x}_{i}^{+})$ corresponding to the cluster centered at $x_{\ell}$ to approximate the function values $g_{0}(x_{\ell})$ and $G(x_{\ell})$ for each $\ell\in[1:d]$. In a second step, based on the approximated function values from the first step, we apply RKHS-based interpolation to approximate the entries in the vector- and matrix-valued functions $g_{0}$ and $G$ in the $\sup$-norm. We briefly highlight that the clustering step in particular alleviates the inherent ill-conditioning of kernel-based interpolation tasks, which manifests in the rapidly decaying eigenvalues of the kernel matrix $\mathbf{K}_{\mathcal{X}}$ [40, Theorem 4.16]. This feature may be seen in the estimate of the subsequent Remark 4.5 as the term including the inverse of the kernel matrix may be controlled by the cluster radius $\varepsilon$.
@@
Output: Approximation $\widehat{f}$ as in Equation 25.
@@
Step 1: Clustering. Choose $\mathcal{X}=\{x_{1},\ldots,x_{d}\}\subset\Omega$.
For each $\ell\in[1:d]$:
$\bullet$ Choose $N$ nearest neighbors $\{x_{\ell,1},\ldots,x_{\ell,N}\}\subset\overline{\mathcal{X}}$ of $x_{\ell}$
$\bullet$ For $j\in[1:N]$: set $u_{\ell,j}=\bar{u}_{i}$ if $x_{\ell,j}=\bar{x}_{i}$
$\bullet$ Define
@@
$H_{\ell}^{*}:=\operatorname*{arg\,min}_{[y\;\;\;Y]\in\mathbb{R}^{n\times m+1}}\Big{\|}\begin{bmatrix}x_{\ell,1}^{+}&\dots&x_{\ell,N}^{+}\end{bmatrix}-\begin{bmatrix}y&Y\end{bmatrix}U_{\ell}\Big{\|}_{F}.$ (27)
@@
Set $\begin{bmatrix}\widehat{g}_{0}&\widehat{G}\end{bmatrix}=\widehat{H}$ and define $\widehat{f}$ as in Equation 25.
We briefly comment on the algorithmic choices of Algorithm 1.
Remark 4.2 (Sampling and clustering).
In Algorithm 1, the points $x_{1},\ldots,x_{d}$ can be seen as points on a “macroscopic” scale whereas the triples $(x_{\ell,i},u_{\ell,i},x_{\ell,i}^{+})$, based on the nearest neighbors, could be understood as data on a “microscopic level”. For the sampling of the data triples $(\bar{x}_{i},\bar{u}_{i},\bar{x}_{i}^{+})$, $i\in[1:\bar{d}]$, and the clustering step in Algorithm 1, various strategies could be applied.
A natural choice for the centers $x_{\ell}$, $\ell\in[1:d]$, would be to minimize the fill distance $h_{\mathcal{X}}$ of the macro level $\mathcal{X}$. Clearly, this choice also influences the precision for the approximation step Equation 27, where we approximate the function values at $\mathcal{X}$ by measurements of the flow at the nearest neighbors. Both quantities, i.e., the fill distance $h_{\mathcal{X}}$ and the radii of the clusters with the $N$ nearest neighbors are reflected in the error bound of Theorem 4.3. For fixed $N$, these quantities may be rendered arbitrarily small, if a sufficiently fine resolution by the micro level data set $\overline{\mathcal{X}}$ is provided (e.g. by grid-based or i.i.d. sampling).
@@
Assume that $U_{\ell}\in\mathbb{R}^{(m+1)\times N}$ in Equation 26 has full rank $m+1$ for all $\ell\in[1:d]$. Then, there exist constants $C,h_{0}>0$ such that for any set $\mathcal{X}\subset\Omega$, chosen in Step 1 of Algorithm 1, with $h_{\mathcal{X}}<h_{0}$, the error between the controlled map $f$ and its surrogate $\widehat{f}$ constructed by Algorithm 1 satisfies, for any $(x,u)\in\Omega\times\mathbb{U}$:
$\|f(x,u)-\widehat{f}(x,u)\|_{\infty}\leq CD(x)(1+\|u\|_{1})$
@@
$D(x)=h_{\mathcal{X}}^{k-1/2}$ $\operatorname{dist}(x,\mathcal{X})\max_{p,q}\|H_{pq}\|_{\mathbb{H}}$
$+\sqrt{2N}\big{(}\max_{\ell\in[1:d]}\|U_{\ell}^{\dagger}\|\big{)}(L_{g}+L_{G}R)\Phi_{n,k}^{1/2}(0)\big{(}\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v\big{)}^{1/2}\varepsilon.$
with the maximal cluster size $\varepsilon=\max_{(\ell,j)\in[1:d]\times[1:N]}\|x_{\ell,j}-x_{\ell}\|$, $H=\begin{bmatrix}g_{0}&G\end{bmatrix}$,
$R=\max\{\|u\|_{\infty}:u\in\mathbb{U}\},\quad\text{and}\quad\mathbbm{1}=\{v\in\mathbb{R}^{d}:v_{i}\in\{\pm 1\}\}.$
@@
Let $\alpha\in\mathbb{R}^{d}$ and $f=\alpha^{\top}\mathbf{k}_{\mathcal{X}}\in V_{\mathcal{X}}$. Then, for all $x\in\Omega$,
$|f(x)|^{2}\,\leq\,\mathsf{k}(x,x)\cdot\alpha^{\top}\mathbf{K}_{\mathcal{X}}\alpha.$
@@
Therefore, the claim follows from $|f(x)|^{2}=|\langle f,\phi_{x}\rangle_{\mathbb{H}}|^{2}\leq\|\phi_{x}\|_{\mathbb{H}}^{2}\|f\|_{\mathbb{H}}^{2}$ and $\|\phi_{x}\|_{\mathbb{H}}^{2}=\langle\phi_{x},\phi_{x}\rangle=\phi_{x}(x)=\mathsf{k}(x,x)$. ∎
@@
Let $\ell\in[1:d]$. First of all, we show that
$|H_{pq}(x_{\ell})-(H_{\ell}^{*})_{pq}|\,\leq\,\sqrt{2N}(L_{g}+L_{G}R)\|U_{\ell}^{\dagger}\|\cdot\varepsilon_{\ell},$ (29)
where $\varepsilon_{\ell}=\max_{j\in[1:N]}\|x_{\ell,j}-x_{\ell}\|$. For this, note that $H_{\ell}^{*}=X_{\ell}^{+}U_{\ell}^{\dagger}$, where $X_{\ell}^{+}:=\begin{bmatrix}x_{\ell,1}^{+}&\dots&x_{\ell,N}^{+}\end{bmatrix}$. Hence, indeed,
@@
$\leq\|e_{p}^{\top}H(x_{\ell})U_{\ell}-e_{p}^{\top}X_{\ell}^{+}\|^{2}\cdot\|U_{\ell}^{\dagger}e_{q}\|^{2}$
$\leq\|U_{\ell}^{\dagger}\|^{2}\sum\nolimits_{s=1}^{N}\Big{|}e_{p}^{\top}H(x_{\ell})Ue_{s}-e_{p}^{\top}X_{\ell}^{+}e_{s}\Big{|}^{2}$
$=\|U_{\ell}^{\dagger}\|^{2}\sum\nolimits_{s=1}^{N}\big{|}e_{p}^{\top}\begin{bmatrix}g(x_{\ell})-g(x_{\ell,s})&G(x_{\ell})-G(x_{\ell,s})\end{bmatrix}\begin{bmatrix}1\\ u_{\ell,s}\end{bmatrix}\big{|}^{2}$
$\leq 2\|U_{\ell}^{\dagger}\|^{2}\sum\nolimits_{s=1}^{N}\Big{(}\|g(x_{\ell})-g(x_{\ell,s})\|^{2}+\|[G(x_{\ell})-G(x_{\ell,s})]u_{\ell,s}\|^{2}\Big{)}$
$\leq 2\|U_{\ell}^{\dagger}\|^{2}N(L_{g}^{2}+L_{G}^{2}R^{2})\varepsilon_{\ell}^{2},$
and Equation 29 follows. Next, for $x\in\Omega$, we have
$\big{|}H_{pq}(x)-\widehat{H}_{pq}(x)\big{|}\leq\big{|}H_{pq}(x)-P_{\mathcal{X}}H_{pq}(x)\big{|}+\big{|}P_{\mathcal{X}}H_{pq}(x)-\widehat{H}_{pq}(x)\big{|}.$
The first summand can be estimated by Theorem 3.7:
$\big{|}H_{pq}(x)-P_{\mathcal{X}}H_{pq}(x)\big{|}\leq Ch_{\mathcal{X}}^{k-1/2}\operatorname{dist}(x,\mathcal{X})\|H_{pq}\|_{\mathbb{H}}.$
For the second, we estimate

### v3

@@ 4. Koopman approximants for control systems
where $x\in\Omega\subset\mathbb{R}^{n}$ and the control $u$ is contained in a bounded set $\mathbb{U}\subset\mathbb{R}^{m}$, $m\leq n$, with $\operatorname{span}\mathbb{U}=\mathbb{R}^{m}$. Here, we assume that the maps $g_{0}:\Omega\to\mathbb{R}^{n}$ and $G:\Omega\to\mathbb{R}^{n\times m}$ are locally Lipschitz continuous, i.e., there is $L_{g}>0$ and $L_{G}>0$ such that
$\|g_{0}(x)-g_{0}(y)\|\leq L_{g}\|x-y\|\qquad\text{and}\qquad\|G(x)-G(y)\|\leq L_{G}\|x-y\|\qquad\forall\,x,y\in\Omega$
Further, all entries of both $g_{0}$ and $G$ are assumed to be in the RKHS $\mathbb{H}=\mathcal{N}_{\Phi_{n,k}}(\Omega)$ for some fixed $k\in\mathbb{N}$. Note that, in view of eq. 3, $\mathcal{N}_{\Phi_{n,k_{1}}}(\Omega)\subset\mathcal{N}_{\Phi_{n,k_{2}}}(\Omega)$ for $k_{1}\geq k_{2}$.
@@
Often, systems eq. 22 are derived from continuous-time control-affine systems governed by the dynamics
$\dot{x}(t):=\tilde{f}(x(t),u(t)):=\tilde{g}_{0}(x(t))+\sum\nolimits_{i=1}^{m}u_{i}(t)\tilde{g}_{i}(x(t))$ (23)
with locally Lipschitz maps $\tilde{g}_{i}:\Omega\rightarrow\mathbb{R}^{n}$, $i\in[0:m]$. Similarly to remark 2.1, we obtain $x(\Delta t;\bar{x},u)=\bar{x}+\int\nolimits_{0}^{\Delta t}\tilde{g}_{0}(x(s;\bar{x},u))\,\mathrm{d}s+\int_{0}^{\Delta t}\tilde{G}(x(s;\bar{x},u))u(s)\,\mathrm{d}s$ for a time step $\Delta t>0$, where $\tilde{G}$ is defined analogously to $G$ above – again tacitly assuming existence and uniqueness of the solution $x(\cdot;\bar{x},u)$ on $0,\Delta t]$. Then, for constant control function $u$ on the interval $(0,\Delta t)$, a Taylor-series expansion of the solution at $t=0$ yields $x(\Delta t;\bar{x},u)=\bar{x}+\Delta t[\tilde{g}_{0}(\bar{x})+\tilde{G}(\bar{x})u]+\mathcal{O}(\Delta t^{2})$, where we have invoked compactness of $\mathbb{U}$ and Lipschitz continuity of $f(\cdot,u)$ and $\frac{\partial}{\partial x}f(\cdot,u)$ on $\overline{\Omega}$ identifying the (constant) control functions with the respective control value. Thus, we obtain a discrete-time system eq. 22 up to an arbitrarily small error if the time step $\Delta t$ is sufficiently small.
In this section, we propose a novel learning architecture to learn Koopman approximants of the control-affine system eq. 22, that allows for flexible sampling of state-control pairs $(\bar{x}_{i},\bar{u}_{i},\bar{x}_{i}^{+})$ with $\bar{x}_{i}\in\Omega$, $\bar{u}_{i}\in\mathbb{U}$, and $\bar{x}_{i}^{+}=f(\bar{x}_{i},\bar{u}_{i})$, $i\in[1:d]$. The result of our proposed kEDMD-based algorithm are control-affine surrogate dynamics
@@
using regression and interpolation in Wendland native spaces. In theorem 4.2, we prove a uniform bound on the approximation error $\|f(x,u)-\widehat{f}(x,u)\|$, similarly as in corollary 2.8 for the approximants $\widehat{F}$ and $\widehat{F}_{\lambda}$ of system eq. DS. Then, we exemplarily apply the derived proportional bound for feedback stabilization. To be more precise, we prove in Section 4.3 that if a feedback stabilizes the surrogate dynamics eq. 24, the desired set point is stabilized in the original system (and vice versa).
@@ 4.1. Koopman approximants, flexible sampling, and error bounds
We propose a novel learning architecture to generate Koopman approximants for control-affine systems with the following two key features: On the one hand, the algorithm can be used with (almost) arbitrary state-control data allowing for highly flexible sampling including i.i.d., grid-based and trajectory-based data generation. On the other hand, we rigorously show bounds on the full approximation error with explicit convergence rates in the infinite-data limit [^9] At this point, the choice of the sampling scheme becomes crucial: while small fill distances can be readily achieved using grid-based methods, trajectory-based sampling typically relies on ergodicity—a restrictive property when considering control applications. and, foremost, allow for controller design with closed-loop guarantees without imposing restrictive invariance assumptions on the dictionary in EDMD. To the best of our knowledge, this combination is, up to now, unique. Koopman-based approximations of control-affine systems in an RKHS framework with flexible sampling were also presented in [3], where the kernel (and thus also the RKHS) is suitably extended to capture the control dependency. Therein, however, no error bounds were provided. In addition, our novel approach allows to counteract numerical ill-conditioning by decomposing the approximation process in two steps labeled as macro and micro level, see the following paragraph for details.
The method proposed in algorithm 1 consists of two steps. First, we determine $d$ clusters of size $N\geq m+1$. Each cluster corresponds to a center $x_{\ell}\in\Omega$, $\ell\in[1:d]$, and its $N$ nearest neighbors. The chosen centers $x_{\ell}$, $\ell\in[1:d]$, determine the fill distance $h_{\mathcal{X}}$. Then, we use the data triples $(\bar{x}_{i},\bar{u}_{i},\bar{x}_{i}^{+})$ corresponding to the cluster centered at $x_{\ell}$ to approximate the function values $g_{0}(x_{\ell})$ and $G(x_{\ell})$ for each $\ell\in[1:d]$. This step addresses the absence of direct measurements of $g_{0}$ and $G$, as these quantities are only available through the lens of the dynamics via a linear combination scaled with the control values. In a second step, we apply RKHS-based interpolation to approximate the entries in the vector- and matrix-valued functions $g_{0}$ and $G$ in the $\sup$-norm. We briefly highlight that the clustering step in particular alleviates the inherent ill-conditioning of kernel-based interpolation tasks, which manifests in the rapidly decaying eigenvalues of the kernel matrix $\mathbf{K}_{\mathcal{X}}$ [47, Theorem 4.16]. This feature may be seen in the estimate of the subsequent theorem 4.2 as the term including the inverse of the kernel matrix may be controlled by the cluster radius $\varepsilon$.
@@
Output: Approximation $\widehat{f}$ as in eq. 24.
@@
Step 1: Clustering. Choose $\mathcal{X}=\{x_{1},\ldots,x_{d}\}\subset\Omega$. For each $\ell\in[1:d]$:
$\bullet$ Choose $N$ nearest neighbors $\{x_{\ell,1},\ldots,x_{\ell,N}\}\subset\overline{\mathcal{X}}$ of $x_{\ell}$ with associated controls $\{u_{\ell,1},\ldots,u_{\ell,N}\}$ and define
@@
$H_{\ell}^{*}:=\operatorname*{arg\,min}\nolimits_{[y\;\;\;Y]\in\mathbb{R}^{n\times m+1}}\big\|\begin{bmatrix}x_{\ell,1}^{+}&\dots&x_{\ell,N}^{+}\end{bmatrix}-\begin{bmatrix}y&Y\end{bmatrix}U_{\ell}\big\|_{F}.$ (26)
@@
Define $\widehat{g}_{0}\in\mathbb{R}^{n}$, $\widehat{G}\in\mathbb{R}^{n\times m}$ by $(\widehat{g}_{0})_{p}=\widehat{H}_{p,1}$ for $p\in[1:n]$, $(\widehat{G})_{p,q}=\widehat{H}_{p,q+1}$ for $(p,q)\in[1:n]\times[1:m]$ and define $\widehat{f}$ as in eq. 24.
We briefly comment on algorithm 1 w.r.t. sampling and clustering: $x_{1},\ldots,x_{d}$ can be seen as points on a “macroscopic” scale whereas the triples $(x_{\ell,i},u_{\ell,i},x_{\ell,i}^{+})$, based on the nearest neighbors, could be understood as data on a “microscopic level”. For the sampling of the data triples $(\bar{x}_{i},\bar{u}_{i},\bar{x}_{i}^{+})$, $i\in[1:\bar{d}]$, and the clustering step in algorithm 1, various strategies could be applied. A canonical choice for the centers $x_{\ell}$, $\ell\in[1:d]$, is to minimize the fill distance $h_{\mathcal{X}}$, e.g., by uniform gridding or a Padua grid as used in this work. Clearly, this choice also influences the precision for the approximation step eq. 26, where we approximate the function values at $\mathcal{X}$ by measurements of the flow at the nearest neighbors. Both quantities, i.e., the fill distance $h_{\mathcal{X}}$ and the radii of the clusters are reflected in the bound of theorem 4.2. For fixed $N$, these quantities may be rendered arbitrarily small, if a sufficiently fine resolution by the micro level data set $\overline{\mathcal{X}}$ is provided (e.g. by grid-based or i.i.d. sampling).
@@
Let $k\geq 1+\frac{(-1)^{n}+1}{2}$ be the smoothness degree of the kernel and $U_{\ell}\in\mathbb{R}^{(m+1)\times N}$ in eq. 25 of full rank $m+1$, $\ell\in[1:d]$. Then, there are constants $C,h_{0}>0$ such that for any $\mathcal{X}\subset\Omega$ chosen in algorithm 1 with $h_{\mathcal{X}}<h_{0}$, the error between the controlled map $f$ and its surrogate $\widehat{f}$ constructed by algorithm 1 satisfies
$\|f(x,u)-\widehat{f}(x,u)\|_{\infty}\leq CD(x)(1+\|u\|_{1})\qquad\forall\,(x,u)\in\Omega\times\mathbb{U},$
@@
$D(x)$ $=\sqrt{2N}\big(\max\nolimits_{\ell\in[1:d]}\|U_{\ell}^{\dagger}\|\big)(L_{g}+L_{G}R)\Phi_{n,k}^{1/2}(0)\big(\max\nolimits_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v\big)^{1/2}\varepsilon$
$\hskip 85.35826pt+h_{\mathcal{X}}^{k-1/2}\operatorname{dist}(x,\mathcal{X})\max\nolimits_{p,q}\|H_{pq}\|_{\mathbb{H}}$
with $H=[g_{0}\ G]$, $R=\max\{\|u\|_{\infty}:u\in\mathbb{U}\}$, maximal cluster size $\varepsilon=\max_{(\ell,j)\in[1:d]\times[1:N]}\|x_{\ell,j}-x_{\ell}\|$, and $\mathbbm{1}=\{v\in\mathbb{R}^{d}:v_{i}\in\{\pm 1\}\}$.
@@
For $\alpha\in\mathbb{R}^{d}$, $f=\alpha^{\top}\mathbf{k}_{\mathcal{X}}\in V_{\mathcal{X}}$, $|f(x)|^{2}\,\leq\,\mathsf{k}(x,x)\cdot\alpha^{\top}\mathbf{K}_{\mathcal{X}}\alpha$ $\forall\,x\in\Omega$.
@@
Consequently, the claim follows from $|f(x)|^{2}=|\langle f,\phi_{x}\rangle_{\mathbb{H}}|^{2}\leq\|\phi_{x}\|_{\mathbb{H}}^{2}\|f\|_{\mathbb{H}}^{2}$ and $\|\phi_{x}\|_{\mathbb{H}}^{2}=\langle\phi_{x},\phi_{x}\rangle=\phi_{x}(x)=\mathsf{k}(x,x)$. ∎
@@
Let $\ell\in[1:d]$. First, we show that
$|H_{pq}(x_{\ell})-(H_{\ell}^{*})_{pq}|\,\leq\,\sqrt{2N}(L_{g}+L_{G}R)\|U_{\ell}^{\dagger}\|\cdot\varepsilon_{\ell},\qquad\varepsilon_{\ell}=\max_{j\in[1:N]}\|x_{\ell,j}-x_{\ell}\|.$ (28)
For this, we have $H_{\ell}^{*}=X_{\ell}^{+}U_{\ell}^{\dagger}$, where $X_{\ell}^{+}:=[x_{\ell,1}^{+}\ \cdots\ x_{\ell,N}^{+}]$. Hence, since $U_{\ell}$ has full row rank, $U_{\ell}U_{\ell}^{\dagger}=I$ such that
@@
$\leq$ $\|e_{p}^{\top}H(x_{\ell})U_{\ell}-e_{p}^{\top}X_{\ell}^{+}\|^{2}\cdot\|U_{\ell}^{\dagger}e_{q}\|^{2}\leq\|U_{\ell}^{\dagger}\|^{2}\sum\nolimits_{s=1}^{N}|e_{p}^{\top}H(x_{\ell})Ue_{s}-e_{p}^{\top}X_{\ell}^{+}e_{s}|^{2}$
$=$ $\|U_{\ell}^{\dagger}\|^{2}\sum\nolimits_{s=1}^{N}\big|e_{p}^{\top}\begin{bmatrix}g_{0}(x_{\ell})-g_{0}(x_{\ell,s})&G(x_{\ell})-G(x_{\ell,s})\end{bmatrix}\begin{bmatrix}1\\ u_{\ell,s}\end{bmatrix}\big|^{2}$
$\leq$ $2\|U_{\ell}^{\dagger}\|^{2}\sum\nolimits_{s=1}^{N}\Big(\|g_{0}(x_{\ell})-g_{0}(x_{\ell,s})\|^{2}+\|[G(x_{\ell})-G(x_{\ell,s})]u_{\ell,s}\|^{2}\Big)$
$\leq$ $2\|U_{\ell}^{\dagger}\|^{2}N(L_{g}^{2}+L_{G}^{2}R^{2})\varepsilon_{\ell}^{2},$
and eq. 28 follows. Next, for $x\in\Omega$, we have
$|H_{pq}(x)-\widehat{H}_{pq}(x)|\leq|H_{pq}(x)-P_{\mathcal{X}}H_{pq}(x)|+|P_{\mathcal{X}}H_{pq}(x)-\widehat{H}_{pq}(x)|.$
We estimate the first summand by $\big|H_{pq}(x)-P_{\mathcal{X}}H_{pq}(x)\big|\leq Ch_{\mathcal{X}}^{k-1/2}\operatorname{dist}(x,\mathcal{X})\|H_{pq}\|_{\mathbb{H}}.$ using theorem 3.7. For the second, we estimate